# rEGGression: an Interactive and Agnostic Tool for the Exploration of Symbolic Regression Models

### v1

@@ 4.1. Available commands
  • top: returns the top $N$ expressions following a user defined criterion.
  • report: reports additional details about the expression with a given id.
@@
  • optimize: (re)optimizes a certain expression by minimizing the specified log-likelihood.
  • insert: inserts and evaluates a new expression, it will display the information about the expression if it already exists in the database.
@@
  • count-pattern: counts the number of occurrences of a certain building block.
@@
  • save: saves the current state of the e-graph.
  • load: loads an e-graph file.
@@
In the following, all these commands will be demonstrated using the nasa_battery_1 dataset. All the files, including the compiled binaries for Linux, MacOS, and Windows are available at https://github.com/folivetti/srtree. The program requires that NLOpt library (Johnson and Schueller, 2021) is correctly installed and accessible in the current system. To start the software, we run the following command line in the console application:
reggression -d nasa_battery_1_10min.csv \
  --load-from nasa_battery.egraph  \
  --distribution MSE
@@ 4.2. Showing the top matching expressions
where [ARG] represents optional arguments and means that it arguments of that type can be repeated. For example, running the command top 3 with size ¡ 5, will result in:
@@
The first column is the id of the expression that can be referred to in other commands, the second column is the expression itself with the variables represented by $x_{i}$ and adjustable parameters as $t_{i}$. The third column is the fitness (assuming maximization) assigned to that expression by the algorithm, the fourth column is the values of the parameters, followed by the number of nodes and the DL, when available. The DL can be calculated when starting the program with the flag –calculate-dl. Alternatively, this can be calculated on-demand with the report command (demonstrated on a later section). The filter of the top command can be composed by adding new with directives, such as:
@@
  • top 3 with cost ¡ 10: will retrieve the top $3$ constrained by the default cost function that assigns a value to each node ($1$ for constants, variables, and parameters, $2$ for binary operators and $3$ for unary operators) and calculates the sum. In a future version, this cost function will be customizable.
@@ 4.4. Disassembling the expression tree
Additionally, by observing the subtrees of two expressions, we may want to test a combination of those to test a new expression our own. This can be done with the insert command.
@@ 4.7. Pareto front
The command pareto (by fitness— by dl) will display the Pareto front where the first objective is either the fitness value or the description length, and the second objective is the size of the expression. To make the retrieval of the front efficient, we exploit the fact that the second objective, size, is a discrete objective with values usually in the range $[1,100]$. In the implementation, we created an array with the size equal to the maximum observed size and each element points to two sorted lists, one with the expressions of that particular size sorted by fitness and the other sorted by DL. Since we only store the e-class ids as integers, this can be made memory efficient allowing a fast retrieval of the Pareto front and subsequent fronts.
@@ 4.9. Distribution of building blocks
For example, when issuing the command distribution with size ¡= 7 limited at 25 by fitness with at least 1000 from top 10000, we obtain Table 1 as a result, listing the top $25$ patterns with size less than or equal to $7$, sorted by the average fitness with a frequency of at least 1 00010001\,0001 000 extracted from the top-10 0001000010\,00010 000 expressions. We should notice that the extraction of building blocks is independent of the equality saturation procedure, thus it will count patterns such as $v_{0}+(v_{1}+v_{2})$ and $(v_{0}+v_{1})+v_{2}$ as two distinct patterns.
@@
Notice that it is important to establish a limit on the pattern size and number of expressions because of the explosive nature of the possible patterns for each expression. All the building blocks of an expression can be extracted with a recursive procedure. A leaf node will contain two building blocks, a pattern variable $v$, that matches everything, and the node itself (e.g., $x$). An unary node will generate the building blocks $v$ and all the building blocks extracted from its child after applying this unary operator. A binary node will generate the build block $v$ together with all the building blocks created after applying the operator to the combinations of building blocks of the left and right children. Considering any expression represented as a complete binary tree, the number of building blocks can be calculated as:
@@
$f(n)$ $=1+f(n-1)f(n-1)$
where $n$ is the number of levels of that tree. Table 2 For trees with up to $5$ levels ($32$ nodes).
@@ 4.10. Saving, loading, and importing
It is also possible to import expressions generated by different symbolic regression algorithms with the command import filename True — False that will import the expressions stored in the file filename. The boolean value indicates whether the parser needs to extract the parameter values from the expression (True) or if they are already provided in the file (False). The extension of the filename will tell the parser which algorithm generated the expressions. For example, for the filename egraph.operon, it will be assumed that the expressions follow the same format as Operon (Burlacu et al., 2020). The supported algorithms are:
Operon (Burlacu et al., 2020)::
an efficient implementation of genetic programming for symbolic regression in C++ with state-of-the-art results (de Franca et al., 2024).
Bingo (Randall et al., 2022)::
a Python implementation of symbolic regression that represents the expressions as a directed acyclic graph with simplification and gradient-based local optimization of parameters.
TIR (de França, 2022) and ITEA (de Franca and de Lima, 2021)::
Haskell implementation of symbolic regression algoithms that constrain the search space to a simplified form.
GOMEA (Virgolin et al., 2017) and SBP (Virgolin et al., 2021)::
applies crossover between interdependent blocks leading to more compact models. SBP uses recursive function inversion to estimate the best modifications to the current model.
EPLEX (La Cava et al., 2016) and FEAT (La Cava et al., 2018)::
genetic programming with $\epsilon$-lexicase selection using semantics to increase the reward of models that perform particularly well in difficult regions, while FEAT incorporates optimizes a linear combination of a set of evolved features using the same selection strategy as EPLEX.
PySR (Cranmer, 2023)::
performs genetic programming with multi-objective optimization on the fitness and complexity while using an island model to improve diversity. The main highlight of this implementation is the possibility to integrate custom operators, custom fitness functions, and semantic constraints.
HeuristicLab (HL) (Wagner and Affenzeller, 2005)::
a graphical interface supporting many optimization and search tasks including symbolic regression.
@@
The imported expressions are inserted into the current e-graph, effectively increasing the number of expressions. This can be used to combine the results from different runs of the same or distinct algorithms. We should notice that many other implementations follows at least one of the formats used by one of these algorithms in the list. More algorithms will be supported on demand. The file to be imported must be a comma separated value with the expression in the first column, a semicolon separated list of coefficient values in the second column, and the fitness value in the third column.
@@ 5. Conclusions
In this paper we present rEGGression, a tool with the main objective to allow the exploration of symbolic regression models generated by multiple sources. The main idea of this tool is to investigate a large database of SR models to explore different alternatives than those selected by the algorithm. The tool brings a text-based interface with the user implementing an SQL-like language in which the user can query for top expressions, search for patterns, and get a report of common building blocks.
@@
From this set of commands, the main highlight is the exploration of the building blocks which can provide frequent patterns observed in expressions with good accuracy, and the average fitness associated with those patterns. This can help users to explorer different combinations of such patterns and enabling them to separate the analysis of the behavior of the model on the more abstract part (the building block) and then on the more specific parts of the expression.
There are many possibilities for future versions of this software. One of them is concerned to the amount of information obtained from a single expression using the report command. Besides the information already shown in this tool, we can also display the expression as a function of observed transformed variables, for example the expression $\theta_{0}sin(x^{2})+\theta_{1}e^{-x}$ can be displayed as $\theta_{0}z_{0}+\theta_{1}z_{1}$, which allows us to treat this as a linear model. Besides that, we could also calculate and display the confidence interval of the parameters, correlation between parameters.

### v2

@@ 4.1. Available commands
  • top: returns the top $N$ expressions following a criterion.
  • report: reports details about the expression with a given id.
@@
  • optimize: (re)optimizes a certain expression by minimizing the specified loss function.
  • insert: inserts, evaluates and displays the information of a new expression.
@@
  • count-pattern: counts the occurrences of a building block.
@@
  • save, load: saves or loads the current state of the e-graph.
@@
In the following, all these commands will be demonstrated using the nasa_battery_1 dataset containing observations of the current state of a battery and its current capacity. The e-graph was generated with eggp(de Franca and Kronberger, 2025) algorithm that makes use of e-graph internally. The hyper-parameter values were: MSE loss function, $200$ generations, population size of $100$, +,,×,÷$+,-,\times,\div,\sin,e,\log,|.|,\operatorname{pow}$ non-terminals set, maximum size of the expression of $20$.
@@ 4.2. Showing the top matching expressions
where [ARG] represents optional arguments and means that that argument can be repeated. For example, running the command top 3 with size ¡ 5, will result in:
@@
The first column is the id of the expression that can be referred to in other commands, the second column is the expression itself with the variables represented by $x_{i}$ and adjustable parameters as $t_{i}$. The third column is the fitness (assuming maximization) assigned to that expression by the algorithm, the fourth column is the values of the parameters, followed by the number of nodes and the DL, when available. The DL can be calculated when starting the program with the flag –calculate-dl. Alternatively, this can be calculated on-demand with the report command (demonstrated on a later section). The filter argument can be composed by adding new with directives, such as:
@@
  • top 3 with cost ¡ 10: will retrieve the top $3$ constrained by the default cost function. The cost function adds up an assigned values to each node ($1$ for constants, variables, and parameters, $2$ for binary operators and $3$ for unary operators). In a future version, the user will be able to customize this cost function.
@@ 4.7. Pareto front
The command pareto (by fitness— by dl) will display the Pareto front where the first objective is either the fitness value or the description length, and the second objective is the size of the expression. The output format is the same of the command top. To make the retrieval of the front efficient, we exploit the fact that the second objective, size, is a discrete objective with limited values. In the implementation, we created an array with the size equal to the maximum observed size and each element points to two sorted lists, one with the expressions of that particular size sorted by fitness and the other sorted by DL. Since we only store the e-class ids as integers, this can be made memory efficient allowing a fast retrieval of the Pareto front and subsequent fronts.
@@ 4.9. Distribution of building blocks
For example, when issuing the command distribution with size ¡= 7 limited at 25 by fitness with at least 1000 from top 10000, we obtain Table 1 as a result, listing the top $25$ patterns with size less than or equal to $7$, sorted by the average fitness with a frequency of at least 1 00010001\,0001 000 extracted from the top-$10000$ expressions. We should notice that the extraction of building blocks is independent of the equality saturation procedure, thus it will count patterns such as $v_{0}+(v_{1}+v_{2})$ and $(v_{0}+v_{1})+v_{2}$ as two distinct patterns.
@@
Notice that it is important to establish a limit on the pattern size and number of expressions because of the fast growing nature of the possible patterns for each expression. All the building blocks of an expression can be extracted with a recursive procedure. A leaf node will contain two building blocks, a pattern variable $v$, that matches everything, and the node itself (e.g., $x$). An unary node will generate the building blocks $v$ and all the building blocks extracted from its child after applying this unary operator. A binary node will generate the build block $v$ together with all the building blocks created after applying the operator to the combinations of building blocks of the left and right children. Considering any expression represented as a complete binary tree, the number of building blocks can be calculated as:
@@
$f(n)$ $=1+f(n-1)^{2}$
where $n$ is the number of levels of that tree. Table 2 shows the number of building blocks for trees with up to $5$ levels ($32$ nodes).
@@ 4.10. Saving, loading, and importing
It is also possible to import expressions generated by different symbolic regression algorithms with the command import filename True — False that will import the expressions stored in the file filename. The boolean value indicates whether the parser needs to extract the parameter values from the expression (True) or if they are already provided in the file (False). The extension of the filename will tell the parser which algorithm generated the expressions. For example, for the filename egraph.operon, it will be assumed that the expressions follow the same format as Operon (Burlacu et al., 2020). The supported algorithms are: Operon (Burlacu et al., 2020), Bingo (Randall et al., 2022), TIR (de França, 2022), ITEA (de Franca and de Lima, 2021), GOMEA (Virgolin et al., 2017), SBP (Virgolin et al., 2021), EPLEX (La Cava et al., 2016), FEAT (La Cava et al., 2018), PySR (Cranmer, 2023), HeuristicLab (HL) (Wagner and Affenzeller, 2005).
@@
The imported expressions are inserted into the current e-graph, effectively increasing the number of expressions. This can be used to combine the results from different runs of the same or distinct algorithms. We should notice that many other implementations follows at least one of the formats used by one of these algorithms in the list. More algorithms will be supported on demand. The imported file must follow a comma separated value format with the expression in the first column, a semicolon separated list of coefficient values in the second column, and the fitness value in the third column. For example: x0^p0 + p1*x1,0.2;3.1,0.89.
4.11. Availability
The source files and compiled binaries for Linux, MacOS, and Windows are available at the srtree symbolic regression library [^5] https://github.com/folivetti/srtree/ repository. The program requires that the NLOpt library (Johnson and Schueller, 2021) is correctly installed and accessible in the current system. The srtree library implements the e-graph data structure and equality saturation algorithms with an extensive set of support functions that can make it easy to extend rEGGression with new functionalities.
4.12. Advantages and limitations
The rEGGression tool is not meant to be a replacement to the other available tools, like the ones mentioned in Sec.2. This work is meant to offer the exploration of a large selection of regression models beyond the common tradeoff between accuracy and complexity, allowing the user to refine the search towards a model that is compatible with the expectations. The most distinct features of rEGGression, when compared to other tools are: the easy-to-use pattern matching of expressions and the reporting of common building blocks. With the pattern matching capability, the user can refine the current selection of top models by enforcing or prohibiting certain building blocks. Likewise, reporting the common building blocks reveals the recurring patterns that may be responsible for the high accuracy of the top models, this can bring additional insights for explaining the behavior of the observations. Finally, in future versions, we will store additional information about the expressions such as whether they are positive, monotonic, and their units when this information is available. This information could be used as additional constraints to the selection of top expressions.
As the main limitations of rEGGression, first of all it is limited to a command line interface. Although the commands are very straightforward, as we add new functionalities it can become increasingly difficult to use. Second, although it supports symbolic expressions as exported by many SR algorithms, it is not possible to integrate this tool into a programmable pipeline (e.g., a Python script that process the data, run the algorithm, and perform post-analysis with rEGGression). Finally, some of the reports could be presented in the form of plots, such as the Pareto front.
@@ 5. Conclusions
In this paper we present rEGGression, a tool with the main objective of exploring symbolic regression models generated by multiple sources. The main idea of this tool is to investigate a large database of SR models to explore different alternatives than those selected by the search algorithm. The tool brings a text-based interface supporting an SQL-like language in which the user can query the top expressions, search for patterns, and get a report of common building blocks.
@@
From this set of commands, the main highlight is the exploration of the building blocks which can provide frequent patterns observed in expressions with good accuracy, and the average fitness associated with those patterns. This can help users interact with the database of expressions, allowing them to test new hypotheses combining accurate expressions or exploiting the building blocks.
There are many possibilities for future versions of this software. One of them is concerned to the amount of information obtained from a single expression using the report command. Besides the information already shown in this tool, we can also display the expression as a function of observed transformed variables, for example the expression $\theta_{0}sin(x^{2})+\theta_{1}e^{-x}$ can be displayed as $\theta_{0}z_{0}+\theta_{1}z_{1}$, which allows us to treat this as a linear model. Besides that, we could also calculate and display the confidence interval of the parameters, and make single point predictions with the confidence intervals.We will also add a simplify command that will run equality saturation on a certain expression to try to reduce its code length.